Let be any positive integer. Use sandwich theorem for sequences to prove that the sequence { 1 n} converges to 0.


1
Expert's answer
2021-01-27T04:06:51-0500

We know that "\{a_n\}" converges to L, or, "\lim_{n \rightarrow \infty}=L" if "\forall" "\epsilon" ">0" "\exist" N"\in" "Z^+" such that "\forall" "n>N",

"|a_n-L|<\epsilon" .

Consider "|\dfrac{1}{ n}-0|=|\dfrac{1}{ n}|=\dfrac{1}{ n}<\epsilon"

"\Rightarrow 1<\epsilon\cdot n \Rightarrow \dfrac{1}{\epsilon}< n \Rightarrow n>\dfrac{1}{\epsilon}"

Proof: Let "\epsilon" ">0" . Choose N">\dfrac{1}{\epsilon}" . Then, "\forall" "n>N""|\dfrac{1}{ n}-0|=|\dfrac{1}{ n}|"

Also, "n>N>\dfrac{1}{\epsilon} \Rightarrow \epsilon>\dfrac1 n \Rightarrow {\dfrac1n}<\epsilon \Rightarrow {\dfrac1{ n}}<\epsilon"

Thus, "|\dfrac{1}{ n}-0|=\dfrac{1}{ n}<\epsilon"

Hence, by above statement, "\{\dfrac{1}{n}\}" converges to 0.